OpenStudy (anonymous):

One of the solutions of x^2+8x=-12 is?

jimthompson5910 (jim_thompson5910):

x^2+8x=-12 x^2+8x+12 = 0 (x+6)(x+2) = 0 I'll let you finish

OpenStudy (anonymous):

x=-6 and -2 :D thank you! That was so simple ._.